OpenStudy (anonymous):

please workout 8t-r=12t

OpenStudy (anonymous):

t=-r/4

OpenStudy (anonymous):

okay how did you get it i don,t understand

OpenStudy (anonymous):

subtract 8t from both sides 8t-8t-r=12t-8t -r=4t then divide both sides by 4 -r/4=4t/4 -r/s=t

OpenStudy (anonymous):

okay so you got rid of the the 8t by subtracting and you divide and where you got the s from

OpenStudy (anonymous):

thank you sp much i was doin it the rite way but just stopped thanks soooo much you don't know how much i mean it omg thanks can u fan me